Use the given conditions to write an equation for the lone in point-slope form slope=6/7, passing through (7,3)
Found 2 solutions by rfer, nerdybill:
Answer by rfer(16322)

y-3=6/7(x-7)

Answer by nerdybill(7384)

Use the given conditions to write an equation for the lone in point-slope form slope=6/7, passing through (7,3)
.
"Point-slope" form is:
y - y1 = m(x - x1)
plugging in the given information:
y - 3 = (6/7)(x - 7)